Atsie (atsie):

I am told to write the linear equation "3x = 5y" in standard form. Standard form is --> Ax + By = C I still don't get the whole idea of standard form and overall it seems plain stupid to me (excuse my bluntness) but what do I do to figure this out properly? What is the purpose of standard form?

Atsie (atsie):

@just_one_last_goodbye @ganeshie8 @Elsa213 Could either one of you be so pleasant to help me please? I'd really appreciate it! ^.^

ganeshie8 (ganeshie8):

What exactly is your question, usefulness of standard form?

Atsie (atsie):

Yes, that is ONE of my questions. What is the purpose of standard form? How do I use it? And also how do I answer the linear equation that I'm given? 3x = 5y ?

ganeshie8 (ganeshie8):

I'll answer the usefulness part shortly. But first could you let me understand how you're stuck on putting the given equation in standard form ?

OpenStudy (er.mohd.amir):

just transfer the terms in one side

Atsie (atsie):

@ganeshie8 Well, I suppose I simply just don't understand the proper way to use it with the equation I'm given.

OpenStudy (er.mohd.amir):

3x+(-5)y=0 or 3x-5y=0 is std form

Atsie (atsie):

@Er.Mohd.AMIR Okay, but why is it standard form? I don't understand.....what is putting it in standard form?

OpenStudy (mathstudent55):

Math is an invention of people. Someone invented standard form for a linear equation to be ax + by = c Other people agreed and decided to use that definition of standard form.

OpenStudy (er.mohd.amir):

since from std form u can derive all other forms live slope-intercept form , intercept form and other also so we first find std form and ten use it for solving problem.

Atsie (atsie):

@mathstudent55 Haha, it IS indeed an invention of people! @Er.Mohd.AMIR Thanks :)

OpenStudy (mathstudent55):

Also, in standard form, a, b, and c should be integers, and a should be non-negative, and there should be no common factors of a, b, and c other than 1. For example, see: http://courses.wccnet.edu/~palay/precalc/22mt01.htm

OpenStudy (mathstudent55):

All things we do in math are conventions and definitions that mathematicians established and were accepted by others. For example, the order of operations is an accepted convention. For example, when we deal with exponents, we know that x^m, where m is an integer greater than or equal to 2, means multiply the base x by itself, using m as the number of factors. That means x^2 = x * x, x^3 = x * x * x, etc. If that is the case then what can x^0 or x^1 possibly mean? There are not enough factors of x to make a multiplication when the exponent is 0 or 1. Well, buy convention, it was agreed that that the definition of x^1 = x for all x, and the definition of x^0 = 1 for all x except x = 0. Once again, you see how we use established conventions and definitions.

OpenStudy (phi):

linear algebra uses standard form: it's the form that "turns into" matrices

OpenStudy (phi):

to do your problem 3x = 5y add -5y to both sides: 3x-5y = 5y-5y on the right side (of the = ) , something minus itself is 0. you get 3x - 5y = 0 or 3x + -5y = 0 (though we don't usually bother to write it this way... it has one more sign, and people usually choose "simple" i.e. 3x-5y=0)

Atsie (atsie):

@phi Thank you so much sir. That really helps out a ton. ^.^